# Where Are We? Evaluating LLM Performance on African Languages

### v1

@@ 5.2 Results and Discussion
We evaluate 13 LMs across four task clusters, maintaining a consistent assessment approach while varying evaluation metrics. Depending on the task, we employ F1, Accuracy, Bleu, and RougeL scores.
As shown in Table 2, in the classification tasks, the closed models Claude-3.5-Sonnet and GPT-4o consistently led in performance, followed by larger LLMs such as Llama3.1-70B and CommandR-Plus-104B. In the language identification task, most models struggled to grasp the task effectively, with only the larger LLMs demonstrating some understanding, suggesting that model scale is crucial for achieving higher accuracy across various multilingual tasks. In the generation task cluster, the larger models performed as expected, with Claude-3.5 and CommandR-Plus leading. There was a noticeable performance increase across all models in the paraphrase task, while the summarization task exhibited a more balanced performance among all models. Considering their size, Llama3.2-3B and Llama3.1-8B performed well within this task cluster. In the MCCR tasks, Llama3.1-70B performed competitively against the closed Claude-3.5 model. In the MGSM task, which involves solving mathematical word problems, most open-source models performed poorly, while Claude-3.5 excelled, achieving an exact match score of 22.34%. In token-level tasks, GPT-4o and CommandR+ consistently delivered strong performances, whereas the Aya-23 model family struggled to grasp the tasks.
The evaluation indicates a landscape of competitive and specialized performance among multilingual and African LLMs, with certain models standing out in some of the tasks. In few-shot evaluation, the closed models followed by Llama3.1-70B and Command-R-Plus, performed significantly better than the other models across multiple tasks. Given their size, the Llama models, especially Llama3.1-8B, perform better than the other models. Additionally, translating into English proved easier than translating from English, reflecting the impact of language representation and data availability on model efficacy (especially at the target side).
Figure 2: Distribution of languages across different clusters. Each bar represents the number of datasets a language appears in, categorized into the clusters, with some languages appearing in all four clusters while others are only present in one or two.
@@
In a limited resource setting, for mathematical QA tasks that fall in the MCCR cluster, Phi3.5 Abdin et al. (2024) ranks top amongst 4 other SLMs for 16 African Languages. In the LLMs category, Gemma 27B has better performance on a higher number of African languages. For reading comprehension tasks, Llama 3.2-3B Dubey et al. (2024) is most suitable in constrained compute resource settings while in scenarios without limits in compute resources, Llama 3.1-70B gives better performance.
Which African language do LMs perform best on?
Most of lama3 models perform best on Swahili. It ranks first in two models and second in one of three models tested. In the Gemma2 models, Aya-x and CommandR+, Swahili and Afrikaans are generally well known. We believe this is due to the availability of data for both languages, with both languages being among the highest resourced languages in Africa.
@@ 6 Data-Performance-Policy Analysis
Figure 3: Distribution of best model performance (Claude 3.5-Sonnet) on different task clusters across African languages in downstream data.
Data Availability and Diversity.
As shown in Figure 2, among the $517$ languages in our benchmark, only $45$ have more than one dataset, while the majority are represented solely by language identification data. Amharic leads with $11$ unique datasets across all clusters, followed by Yorùbá with ten and Hausa with nine. A total of six languages have only one dataset beyond language identification, while the remaining $36$ have between two and eight datasets each. It is important to note that the disparity in dataset availability is closely linked to language policy. The $45$ languages with multiple datasets generally hold official status or are spoken by a significant majority in their respective countries. Government policies that designate certain languages as official often yield greater institutional support, driving increased documentation, education, and media presence. For example, Amharic’s official status in Ethiopia contributes to its relatively rich dataset availability, while Yorùbá and Hausa, both widely used in Nigeria for regional governance and media, similarly benefit from enhanced linguistic resources.
Conversely, languages lacking official recognition or widespread institutional support receive minimal investment in digital and linguistic infrastructure. The dominance of a few languages within policy frameworks perpetuates data scarcity for many Indigenous African languages, thereby limiting their inclusion in NLP advancements and reinforcing digital inequalities. Addressing this imbalance requires policy shifts that prioritize multilingual data collection, increased funding for low-resource language research, and the integration of diverse languages into education, media, and digital communication.
Data Quality.
@@
Here, we analyze performance of our best model, Claude 3.5-Sonnet. As shown in Figure 3, the model faces varying degrees of difficulty when handling African languages. Sub-figure 3(a) demonstrates that for generation tasks, even the strongest models struggle to produce high-quality text in African languages, with most scores falling below 15 BLEU/ROUGE. In contrast, sub-figure 3(b) shows that while many languages achieve relatively strong classification performance (over 80% accuracy), a substantial portion still falls into the 60%–80% range, underscoring a clear performance gap. Finally, sub-figures 3(c) and 3(d) reveal that translation into English is notably easier than translation out of English or French into African languages, where BLEU scores drop from an average of around $19$ down to nine or $11$, respectively. This finding suggests that even the best model struggles more with producing fluent, faithful output in African languages than with understanding them—underscoring the urgent need for better data and model development tailored to these languages. Refer to Appendix E for a similar analysis leveraging our empirical language identification evaluation.
Apart from our empirical analyses, the limited availability and diversity of datasets for African languages in NLP benchmarks has significant implications for developing robust language technologies. Limited Model Performance and Generalization – Models trained on a single dataset per language risk overfitting to specific domains or styles, reducing their ability to generalize across different contexts, which is especially problematic for African languages with high dialectal variation. Bias Toward Well-Resourced Languages – Languages with more datasets naturally perform better in multilingual models, reinforcing disparities and leading to poor performance for underrepresented African languages. Restricted Downstream Applications – Many NLP applications, such as machine translation and question answering, require diverse datasets; if a language is represented only by language identification data, it cannot support the development of more complex language technologies. Challenges in Transfer Learning – Transfer learning techniques, like multilingual fine-tuning, rely on cross-lingual similarities and sufficient data; when African languages suffer from sparse training data, they are unable to effectively leverage knowledge transfer from resource-rich languages.
@@ 8 Conclusion
Acknowledgments
We acknowledge support from Canada Research Chairs (CRC), CLEAR Global, the Natural Sciences and Engineering Research Council of Canada (NSERC; RGPIN-2018-04267), the Social Sciences and Humanities Research Council of Canada (SSHRC; 895-2020-1004; 895-2021-1008), Canadian Foundation for Innovation (CFI; 37771), Digital Research Alliance of Canada, [^2] https://alliancecan.ca and UBC ARC-Sockeye. [^3] https://arc.ubc.ca/ubc-arc-sockeye We also thank Clémentine Fourrier and Nouamane Tazi from HuggingFace for providing feedback on this work. The findings and conclusions contained within this work those of the authors and do not necessarily reflect positions or policies of any supporters.
@@ Appendix B Benchmark
Table B1 shows the comparison between Sahara with others
@@ Appendix D Dataset Description
Table D2 describes all datasets included in Sahara
@@ Appendix E Data-Performance-Policy Analysis
(a) Languages with an F1 score of $50$ or higher.
(b) Languages with an F1 score below $50$.
Figure E2: Overall distribution of Claude-3.5-Sonnet’s performance on language identification.
Figure E2 shows the overall distribution of Claude-3.5-Sonnet’s (best model) performance on language identification. In sub-figure 2(a), the vertical axis represents the F1 score, and each point along the horizontal axis corresponds to a specific African language for which the model achieves an F1 score of $50$ or higher. Only around $50$ languages fall into this category, with their performance ranging from just above $50$ F1 to approximately $90$ F1. Meanwhile, sub-figure 2(b) depicts the majority of African languages, all of which register F1 scores below $50$. These results indicate that while the model manages moderate or better performance on a limited subset of languages, it struggles substantially with the remaining ones. Overall, this visualization underscores that the model’s ability to correctly identify and process most African languages remains limited, highlighting a need for more comprehensive data coverage and targeted model improvements.

### v2

@@ 5.2 Results and Discussion
We evaluate $24$ LMs across four task clusters, maintaining a consistent assessment approach while varying evaluation metrics. Depending on the task, we employ Exact Match, F1, Accuracy, spBLEU1K Elmadany et al. (2024), and RougeL scores.
As shown in Table 2, in the classification tasks, the closed models Claude-4-Sonnet and GPT-4.1 consistently lead in performance, followed by larger LLMs such as Gemma3 (27B) and CommandR-Plus-104B. In the language identification task, most models struggle to grasp the task effectively, with only the larger LLMs demonstrating some understanding, suggesting that model size is crucial for achieving higher accuracy across various multilingual tasks. In the generation task cluster, larger models perform as expected, with GPT-4.1 and Llama-3.1 (70B) achieving the best results. There is a noticeable performance increase across all models in the paraphrase task, while the summarization task exhibits a more balanced performance among all models. Considering their size, Phi-3.5 and Gemma3 (4B) perform well within this task cluster. In the MCCR tasks, Command-A ( [^1] B) performs competitively against the closed Claude-4-Sonnet model. In the MGSM task, which involves solving mathematical word problems, most open-source models perform poorly, while Claude-4-Sonnet excels, achieving an exact match score of $22.34$%. In token-level tasks, GPT-4.1 and Command-A ( [^1] B) consistently deliver strong performances, whereas the Aya model family struggle to grasp the tasks.
Results demonstrate a nuanced landscape of performance across multilingual and African LLMs, with notable variations depending on parameter size and task type. In few-shot settings, the closed model Claude-4-Sonnet achieves the highest overall average score (40.82), outperforming all open models across most task clusters. Among open models, Command-A ( [^1] B) and Gemma-3 (27B) also exhibit strong performance, particularly in downstream tasks such as reasoning tasks, where Gemma2 (27B) reaches top score of 82.49 on Squad-QA and Command-A achieves 61.91 on MMLU task. For smaller parameter regimes, Gemma-2 (2B) and Phi-4 (3.8B) demonstrate the best performance on their model size. Furthermore, Gemma2 (9B, 12B) show consistent mid-to-high performance across various clusters, but do not surpass the larger models in aggregate. Importantly, translating into English is consistently easier for the models than translating from French, likely reflecting disparities in training data distribution and language representation on the target side. These findings highlight not only the dominance of closed LLMs and very large open models in multilingual tasks, but also the critical influence of resource availability and model scale on task-specific outcomes.
@@
In a limited resource setting, for mathematical QA tasks that fall in the MCCR cluster, Phi-4 ranks top among SLMs for 16 African Languages. In the LLMs category, Command-A ( [^1] B) has better performance on a higher number of African languages. For reading comprehension tasks (belebele), Phi-4 is most suitable in constrained compute resource settings while in scenarios without limits in compute resources, Llama 3.3-70B gives better performance.
Which African language(s) do LMs perform best on?
The models achieve their highest performance on a small set of well-resourced African languages—chiefly Hausa, Swahili, Yorùbá, and Afrikaans—across diverse NLP tasks. Both open-source models like Command-A ( [^1] B) and Aya-35B and proprietary models such as Claude-4-Sonnet consistently excel on these languages, reflecting their ample labeled and parallel data. This trend spans simple classification and NER tasks to complex reasoning benchmarks (e.g., MMLU, XLNI) and machine translation, underscoring that data availability, rather than linguistic typology, is the primary driver of model effectiveness.
Models explicitly designed with African linguistic diversity, such as Aya-35B, also achieve strong results on Southern Sotho and Lingala—languages that have benefited from recent, community-led data initiatives. Swahili, in particular, ranks among the top languages for Llama3 models, likely due to its standardized orthography, regular morphology, and extensive bilingual corpora. Afrikaans and Swahili similarly perform well within Gemma2 and Command-R-Plus (104B), supported by their inclusion in massive multilingual datasets (e.g., mC4, OSCAR, WMT) and, in Afrikaans’s case, established government-supported digital media.
Overall, our findings reveal that LM performance correlates strongly with the volume and diversity of available data. Underperformance on low-resource African languages is not a function of inherent complexity but of underrepresentation in training data. Bridging this gap will require concerted efforts to collect, annotate, and standardize datasets that capture the dialectal, sociolinguistic, and orthographic richness of all African languages—an essential step toward improved model equity and real-world utility across the continent’s full linguistic spectrum.
@@ 6 Data-Performance-Policy Analysis
Data Availability and Diversity.
As shown in Figure E3, among the $517$ languages in our benchmark, only $45$ have more than one dataset, while the majority are represented solely by language identification data. Amharic leads with $11$ unique datasets across all clusters, followed by Yorùbá with ten and Hausa with nine. A total of six languages have only one dataset beyond language identification, while the remaining $36$ have between two and eight datasets each. It is important to note that the disparity in dataset availability is closely linked to language policy. The $45$ languages with multiple datasets generally hold official status or are spoken by a significant majority in their respective countries. Government policies that designate certain languages as official often yield greater institutional support, driving increased documentation, education, and media presence. For example, Amharic’s official status in Ethiopia contributes to its relatively rich dataset availability, while Yorùbá and Hausa, both widely used in Nigeria for regional governance and media, similarly benefit from enhanced linguistic resources. Conversely, languages lacking official recognition or widespread institutional support receive minimal investment in digital and linguistic infrastructure. The dominance of a few languages within policy frameworks perpetuates data scarcity for many Indigenous African languages, thereby limiting their inclusion in NLP advancements and reinforcing digital inequalities.
It is noteworthy that availability of NLP data is not necessarily determined by the number of speakers and that other variables including language prestige, literacy rates, and economic incentives can be involved. This becomes evident when comparing speaker populations with actual data availability. For example, Hausa (73 million speakers), Swahili (80 million), and Zulu (28 million) are categorized as hopefuls—languages with some available labeled datasets. In contrast, Naija (Nigerian Pidgin)—despite having approximately 153 million speakers—is classified as a left-behind language Joshi et al. (2020), largely due to decades of stigmatization and negative public attitudes Oyebola and Ugwuanyi (2023), which have significantly contributed to the scarcity of accessible and annotated linguistic resources. Notably, several non-African languages with significantly smaller speaker populations—such as Catalan (5 million), Finnish (10 million), and Swedish (10.5 million)—are regarded as high-resource, primarily as a result of sustained efforts in documentation, digitization, and technological integration. This discrepancy underscores a broader systemic issue in language representation: it is not just linguistic demographics that shape digital inclusion, but also factors such as historical language prestige, policy decisions, digital infrastructure, and the extent of prior documentation.
These intersecting factors ultimately determine whether a language is integrated into AI systems or remains marginalized. Addressing this imbalance requires a shift in policy and practice to prioritize equitable digital representation. This includes targeted investments in multilingual data collection, increased funding for research on historically excluded and low-resource languages, and the intentional integration of these languages into education systems, media platforms, and digital technologies. Such efforts must also consider the sociopolitical and infrastructural barriers that shape language inclusion in AI, ensuring that underrepresented languages are not only documented but actively supported within the broader digital ecosystem.
Data Quality.
@@
Figure 2: Distribution of best model performance (Claude-4-Sonnet) on different task clusters across African languages in downstream data.
@@
Here, we analyze performance of our best model, Claude-4-Sonnet. As shown in Figure 2, the model faces varying degrees of difficulty when handling African languages. Figure 2(a) demonstrates that for generation tasks, even the strongest models struggle to produce high-quality text in African languages, with most scores falling below 15 BLEU/ROUGE. In contrast, Figure 2(b) shows that while many languages achieve relatively strong classification performance (over 80% accuracy), a substantial portion still falls into the 60%–80% range, underscoring a clear performance gap. Finally, Figures 2(c) and 2(d) reveal that translation into English is notably easier than translation out of English or French into African languages, where BLEU scores drop from an average of around $19$ down to nine or $11$, respectively. This finding suggests that even the best model struggles more with producing fluent, faithful output in African languages than with understanding them—underscoring the urgent need for better data and model development tailored to these languages.
Apart from our empirical analyses, the limited availability and diversity of datasets for African languages in NLP benchmarks has significant implications for developing robust language technologies. Limited Model Performance and Generalization – Models trained on a single dataset per language risk overfitting to specific domains or styles, reducing their ability to generalize across different contexts, which is especially problematic for African languages with high dialectal variation. Bias Toward Well-Resourced Languages – Languages with more datasets naturally perform better in multilingual models, reinforcing disparities and leading to poor performance for underrepresented African languages. Restricted Downstream Applications – Many NLP applications, such as machine translation and question answering, require diverse datasets; if a language is represented only by language identification data, it cannot support the development of more complex language technologies. Challenges in Transfer Learning – Transfer learning techniques, like multilingual fine-tuning, rely on cross-lingual similarities and sufficient data; when African languages suffer from sparse training data, they are unable to effectively leverage knowledge transfer from resource-rich languages. We also analyze the performance of Command-A in Figure E2.
@@ 10 Ethical Considerations
Acknowledgments
We acknowledge support from Canada Research Chairs (CRC), CLEAR Global funding from the Gates Foundation, the Natural Sciences and Engineering Research Council of Canada (NSERC; RGPIN-2018-04267), the Social Sciences and Humanities Research Council of Canada (SSHRC; 895-2020-1004; 895-2021-1008), Canadian Foundation for Innovation (CFI; 37771), Digital Research Alliance of Canada, [^5] https://alliancecan.ca and UBC ARC-Sockeye. [^6] https://arc.ubc.ca/ubc-arc-sockeye We also thank Clémentine Fourrier and Nouamane Tazi from HuggingFace for providing feedback on this work. The findings and conclusions contained within this work those of the authors and do not necessarily reflect positions or policies of any supporters.
@@ Appendix B Benchmark
Table B1 shows the comparison between Sahara with other benchmarks.
@@ Appendix D Dataset Description
Table D2 describes all datasets included in Sahara.
@@ Appendix E Data-Performance-Policy Analysis
(a) Generation tasks.
(b) Classification tasks.
(c) MCCR tasks.
(d) MT tasks grouped.
Figure E2: Distribution of best open weight model performance (Command-A) on different task clusters across African languages in downstream data.
Figure E3: Distribution of languages across different clusters. Each bar represents the number of datasets a language appears in, categorized into the clusters, with some languages appearing in all four clusters while others are only present in one or two.
(a) Command-A performers.
(b) Claude-4-Sonnet performance.
Figure E4: Performance of Command-A and Claude-4-Sonnet on token-level task cluster across African languages in downstream data.
SLM LLM Closed LLM
Cluster Identifier Metric Shots S1 S2 S3 S4 S5 S6 L1 L2 L3 L4 L5 L6 L7 L8 L9 L10 L11 L12 L13 L14 L15 L16 CL1 CL2
Classification xlni Acc. 5 33.6 33.7 36.7 41.18 40.4 46.68 48.08 55.06 34.3 40.69 37.09 61.09 57.86 65.58 44.2 38.19 51.89 48.38 52.09 35.87 46.77 59.67 69 69.6
lid $F_{1}$ 10 0 0 0.87 0.26 0.85 0.77 2.82 3.6 0 1.49 2.27 1.95 4.61 4 1.81 3.34 3.26 3.07 3.3 3.05 3.32 4.57 3.73 4.48
news $F_{1}$ 3 0.86 0.35 0.15 17.54 0 6.64 21.94 0 2.06 24.06 0.12 25.83 7.83 35.56 28.03 0.61 35.26 33.31 34.81 3.45 30.46 28.66 40.09 32.19
sentiment $F_{1}$ 5 38.51 24.75 30.45 11.05 28.77 1.99 36.09 42.47 38.4 28.36 36.07 34.56 20.55 46.35 35.86 25.94 24.51 45.82 35.89 19.98 37.07 47.14 55.67 57.52
topic $F_{1}$ 3 8.73 31.62 20.3 19.79 12.46 17.15 48.11 39.35 34.95 28.55 42.12 44.3 31.8 70.73 36.47 43.74 64.89 56.48 48.73 52.09 50.96 53.16 67.93 76.57
Avg. 16.34 18.08 17.69 17.96 16.50 14.65 31.41 28.10 21.94 24.63 23.53 33.55 24.53 44.44 29.27 22.36 35.96 37.41 34.96 22.89 33.72 38.64 47.28 48.07
Generation mt_eng2xx spBleu1K 5 1.79 3.63 3.45 3.25 2.12 3.4 7.07 7.61 1.72 4.37 2.79 6.24 6.91 4.56 3.73 3.29 8.84 8.7 9.36 1.4 6.06 8.22 12.18 12.36
mt_fra2xx spBleu1K 5 0.79 0.36 0.74 1.05 1.25 1.32 0.73 0.95 0.7 1.19 0.66 2.03 1.67 2.05 1.15 0.82 1.93 1.87 1.77 1.05 1.04 2.3 3.41 3.42
mt_xx2xx spBleu1K 5 0.4 0.45 0.42 0.42 0.36 0.18 0.43 0.63 0.72 0.55 0.46 0.71 0.9 1.39 0.55 0.17 1.28 1.05 1.37 0.67 0.38 0.85 5.15 4.44
paraphrase spBleu1K 5 23.2 28.88 31.15 19.12 25.93 17.85 28.73 25.83 31.68 31.33 23.66 17.26 21.34 15.37 32.13 30.68 26.86 19.64 27.79 19.23 36.06 26.06 23.2 20.35
summary rougeL 2 10.03 3.24 4.37 10.07 0.91 13.34 15.1 10.43 1.45 12.74 5.92 16.16 15.73 17.13 13.27 4.19 16.92 17.19 16.4 15.46 14.88 16.09 5.99 16.3
title spBleu1K 2 2.14 2.81 0.25 2.3 0.03 5.78 7.48 5.78 3 5.94 1.58 7.75 7.51 8.65 6.83 3.2 11.07 10.34 8.78 7.32 6.69 8.65 13.62 9.48
Avg. 6.39 6.56 6.73 6.04 5.10 6.98 9.92 8.54 6.55 9.35 5.85 8.36 9.01 8.19 9.61 7.06 11.15 9.80 10.91 7.52 10.85 10.36 10.59 11.06
MCCR mmlu Acc. 5 24.09 28.5 28.68 27.29 28.9 30.99 31.88 40.9 26.4 44.51 22.1 42.09 44.9 60.71 51.5 24 56.51 59.3 50.42 59.81 29.8 61.91 81.6 58.6
mgsm ExactM 5 2.6 3.21 0 1.2 3.39 0.7 2.9 2.8 3.3 2.7 1.5 2.3 5.3 0.6 4.4 3 10.1 7.51 7.6 4.5 4.39 11.1 45.9 29.5
belebele Acc. 5 27.71 14.71 25.2 29.31 31.31 24.69 22.7 26.93 25.09 28.24 26.33 25.63 31.04 32.34 33.64 26.61 30.4 35.92 33.72 30.83 20.99 29.12 36.6 39.7
squad_qa $F_{1}$ 5 57.45 70.89 67.32 43.13 71.31 74.97 78.04 80.91 52.66 59.48 63.41 73.78 82.49 79.16 69.25 64.64 77.65 76.35 66.25 53.92 74.64 80.05 78.02 76.11
Avg. 27.96 29.33 30.30 25.23 33.73 32.84 33.88 37.89 26.86 33.73 28.34 35.95 40.93 43.20 39.70 29.56 43.67 44.77 39.50 37.27 32.46 45.55 60.53 50.98
Tokens ner $F_{1}$ 5 0 0.78 1.76 2.99 0.13 3.65 2.26 7.49 2.35 3.96 0.2 6.03 13.14 11.4 8.51 1.52 6.86 10.54 8.64 6.04 4.3 16.88 35.06 27.04
phrase $F_{1}$ 5 14.79 28.5 29.44 11.52 24.91 27.73 15.95 29.35 29.42 16.55 30.87 13.59 30.01 23.97 26.34 23.09 25.55 23.13 16.78 25.13 29.44 31.05 36.69 36.45
pos $F_{1}$ 5 8.83 6.73 10.56 8.04 10.3 10.46 9.13 13.2 9.81 11.68 12.52 13.86 19 13.99 18.36 11.4 14.13 15.04 11.52 20.76 13.71 27.54 62.84 38.67
Avg. 7.87 12.00 13.92 7.52 11.78 13.95 9.11 16.68 13.86 10.73 14.53 11.16 20.72 16.45 17.74 12.00 15.51 16.24 12.31 17.31 15.82 25.16 44.86 34.05
Overall 14.64 16.49 17.16 14.19 16.78 17.10 21.08 22.80 17.30 19.61 18.06 22.25 23.80 28.07 24.08 17.75 26.57 27.05 24.42 21.25 23.21 29.93 40.82 36.04
Table E4: Few-shot evaluation across different task clusters/tasks. The highest score for each individual task is in bold green, while the best average score across each task cluster (and overall score, calculated as average of task cluster averages) is in bold orange. Underline: refers to the best score across the open source/weights LLMs. Shots: number of shots. SLMs: S1 Llama3.2 (1B), S2 Llama3.2 (3B), S3 Gemma2 (2B), S4 Phi-3.5 (3.8B), S5 Phi-4 (3.8B), and S6 Gemma3 (4B). LLMs: L1 Llama3.1 (8B), L2 Gemma2 (9B), L3 Aya (8B), L4 Babel (9B), L5 Command-R-Plus-R7B (8B), L6 Gemma3 (12B), L7 Gemma2 (27B), L8 Gemma3 (27B), L9 DeepSeek-R1-Distill-Qwen (32B), L10 Aya (35B), L11 Llama3.1 (70B), L12 Llama3.3 (70B), L13 DeepSeek-R1-Distill-Qwen (70B), L14 Babel (83B), L15 Command-R-Plus (104B), L16 Command-A ( [^1] B). CLosed LLMs: CL1 Claude-4-Sonnet, CL2 GPT-4.1.